Dr. Wilson purchased a luxury car for $32,500. This is $7,000 more than twice the price her husband paid for a commuter car earlier that year. Find the price of the commuter car.
Answer by checkley77(12844)

32,500=7,000+2x
32,500-7,000=2x
25,500=2x
x=25,500/2
x=12,750